# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

### v1

@@ 3.3. Example: Learning from dogs, cats, and frogs
Every token in this hierarchy is retokenizable into only previously learned parts. This guarantees that inference can recursively decompose any learned token, and that the vocabulary is self-consistent and locally compositional. For example, given context frog, the model knows to terminate right-inference because no merger of $\tilde{v}^{(4)}_{\bf frog}$ with another token exists as a feature map – defined by non-vanishing action of $P_{n}P_{n-1}\cdots P_{2}$ (e.g Fig 1). Nevertheless, the model can still infer frogs as the best non-smooth construction due to local interactions $g^{(2)}_{\bf og}$ $g^{(3)}_{\bf ogs}$.
@@
This latter subtlety highlights the hybrid sub-symbolic/automaton nature of the model, where tokenization [^2] This dynamic tokenization is distinct from both tokenization (Schuster and Nakajima, 2012; Gage, 1994) & entropy-based approaches to LLM’s (Li et al., 2025; Zhang et al., 2024; Agrawal et al., 2024; Simonds, 2025). LLM’s assume pretokenized input (Phuong and Hutter, 2022; Gage, 1994; Schuster and Nakajima, 2012). plays a dynamical role during inference to determine boundaries, and therefore imbues the model with a notion of accepted string akin to automata. In energetic terms, the accepted strings live in a low-energy manifold (the smooth tokenset), and therefore are the most-likely configurations to be observed given the context. We may therefore think of the automaton as living in the ground state, with its language being the ground state configurations (Reinhart and las Cuevas, 2022). This approximate behavior is made precise by the introduction of projectors (and therefore a consequence of learning higher-order correlations while requiring locality).
@@
Figure 1. Summarizing the relationship between tokenizability, locality, and hierarchy. Three words from a random language (Eugenio, 2025) (shown in blue): yjjfqi, kltilj, & yjjfgsp.
@@ 4. The benefits of thinking about it
The hierarchical short-term memory stores learned words as smooth tokens encoded via a bundle of entangled feature maps ($P_{n}\cdots P_{2}$) and local interactions ($g^{(n)}$). This allows for the model to determine word boundaries (Fig 1), as well as retrieve words from context.
However, this places a number of limitations which limit its scalability. First, it learns all feature maps as entangled, regardless of if they share overlapping graphs – this is highly impractical [^3] Because the feature maps are entangled, the projector size scales as the square of the number of learned features ($n$-grams) (Eugenio, 2025). Because the $P_{n}$ are sparse, a tradeoff can be made to improve memory, however at the cost of losing the representational power of dense matrices. Additionally, memories would have to be stored & accessed from a singular large hierarchy, regardless of whether most of the memory traces are unrelated to the given context. And we noticed the total number of $n$-grams in a real text scales as a power-law (Eugenio, 2025).. Further, word length is finite – stored strings are limited to regular-language-like structure.
The model overcomes these limitations by employing a thinking-like process we call replay relearning. This procedure binds the distributed features of a token sequence (e.g a word object) into a new long-term embedding neuron $a_{\alpha}$, thus disentangling the feature maps into a series of smaller word-specific feature maps. These can later participate in inference independent of the original $g^{(n)}$ (which can be forgotten).
@@ 4.1. The Replay Game (Model Thinking)
[…]
The result is a key-value memory binding (Tyulmankov et al., 2021; Gershman et al., 2025; Ba et al., 2016): the neuron $a_{\alpha}$ becomes a long-term embedding associated with the full replayed word. Over time, this mechanism constructs a library of long-term memory neurons $\{a_{\alpha}\}$, each associated with a unique word and its compositional features. Once learned, these embeddings enable fast inference via pattern completion: a partial word (e.g. a suffix or prefix) can trigger an embedding, which drives next-token prediction.

### v2

@@ 3.3.1. Topology of the Ground State Automaton
Note that, given context frog, the model knows to terminate right-inference because no merger of $\tilde{v}^{(4)}_{\bf frog}$ with another token exists as a feature map – defined by non-vanishing action of $P_{n}P_{n-1}\cdots P_{2}$. Nevertheless, the model can still infer frogs as the best non-smooth construction due to local interactions $g^{(2)}_{\bf og} &$ $g^{(3)}_{\bf ogs}$.
@@
This latter subtlety highlights the hybrid sub-symbolic/automaton nature of the model, where tokenization [^7] This dynamic tokenization is distinct from both tokenization (Schuster&Nakajima2012_SubwordTokenization; Gage1994_BPE) & entropy-based approaches to LLM’s (Li&all2025_EntropyAwareGeneration; Zhang&all2024_EntropyAwareGeneration; Agrawal&all2024_EntropyAwareGeneration; Simonds2025_EntropyAwareGeneration). LLM’s assume pretokenized input (Phuong&Hutter2022review_DeepMind_Transformers; Gage1994_BPE; Schuster&Nakajima2012_SubwordTokenization). plays a dynamical role during inference to determine boundaries, and therefore imbues the model with a notion of accepted string. In energetic terms, the accepted strings live in a low-energy translationally-invariant manifold, and therefore are the most likely configurations to be observed given the context. (Thus in our example, frogs is well-formed but not accepted.)
Their corresponding automata live in the ground state, with their formal language being the ground state configurations (Reinhart&DeLasCuevas2022_IsingGrammar). A projection onto the ground state manifold [^8] This ground-state projection is similar to, for example, how high magnetic fields and low temperatures can effectively project the electronic states of a quantum Hall magnetic onto its lowest Landau level (EugenioDag2020scipost). Model parameters & environmental conditions constrain the model onto a gapped isolated subspace in the spectrum. guarantees an emergent automatonic behavior. Local discorrelations manifest as jumps in energy, up out of the ground state manifold. However, the projection sharpens this behavior to exactness, such that the projectors can be used in placed of energy/entropy jumps. Hence the distinction of accepted vs well-formed as one of being respectively in or just above the manifold.
Further we stress that the ground state is topological. This is a non-trivial statement, not generically guaranteed for every ground state behavior of a physical system, and arises here because a global order parameter is learned in a strictly local way. The topology of the retokenizer is the topology of the DAG graph, with the strict every $n$-gram feature be composed of (n1)(n-1)-gram features smoothness constraint. The terminal nodes of the DAG are special features towards which the autoregressive flows are directed. Thus retrieval of terminal nodes is topologically guaranteed (e.g Fig 3). [^9] This parallels recent industry advances, which exploit bifix-free data structures to reduce interference during retrieval (Augeri2025_HyperTokens; Augeri2008thesis) in LLMs.
Enforcing this constraint is equivalent to demanding gauge invariance: define left (right) retokenization operators
$\mathcal{L}_{n}:\tilde{v}^{(n-1)}v\rightarrow v\tilde{v}^{(n-1)}$
$\mathcal{R}_{n}:v\tilde{v}^{(n-1)}\rightarrow\tilde{v}^{(n-1)}v$
as series of projections & reprojections using $P_{n}$ & $P_{L,n}$. The product $\mathcal{R}\mathcal{L}$ is the projector onto the smooth tokenset [^10] Note the gauge structure of retokenization arises because we use dense matrices (projectors) to describe the information of a sparse graph. However, the brain is not dense, and therefore enforcing gauge invariance doesn’t implicitly mean anything, as gauges are non-physical artifacts of the dense representation. In our companion work (Eugenio2025_HebbianLanguage), we demonstrate how the equivalent DAG constraint is a likely consequence of the combination of translational invariance plus a simple threshold function., and is equivalent to applying a DAG mask. Thus we can interpret the exact retrieval of the global order parameter as being directly tied to the ability of the model to perform both right & left token prediction consistent within a singular mathematical structure. (Or equivalently converging to the terminal nodes of the sparse graph.)
The ground state topology enforces a global compositional word-forming vs non-word forming distinction: you are either in the graph or you’re not. Non-word forming patterns aren’t just local discorrelations, measurable as a jump in entropy, but topological defects tied to the global composition rule. If a pattern is not a word forming pattern, then it necessarily contains a boundary (e.g Fig 2).
Unlike estimating continuous information, where discorrelations reflect changes in a physical medium, what defines a “good” vs “bad” word pattern can be arbitrary. Presenting the model one word at a time during training allows it to learn what a “good” pattern is. Yet human sub-word structure is not arbitrary, but has a specific structure reflecting the underlying neural topology. This becomes apparent when the model is allowed to self-organize (i.e trained on unstructured noise or grown randomly). We discuss this more in Section 4.
@@
Figure 2. Summarizing the relationship between retokenization, locality, and hierarchy. Three words from a random language (Eugenio2025_HebbianLanguage) (shown in blue): yjjfqi, kltilj, & yjjfgsp.
3.3.2. Quick note about spaces
Neither punctuation nor spacing are necessary elements of writing:
$\texttt{youcanreadthisoranysuchtextjustfine}.$
Historically, human writing existed without specialized tokens indicating boundary. This is best illustrated by alphabets, whose fundamental tokens are meaningless beyond their sound associations. Spaces (and other punctuation) are not real tokens inside human language patterns, but are external annotations of statistical boundaries.
In the context of the model, these boundaries are high energy discorrelations. We can lower the energy of a string by inserting a zero vector, e.g $v_{\bf y},v_{\bf o},v_{\bf u},\vec{0},v_{\bf c},v_{\bf a},v_{\bf n}]$. But $\vec{0}$ is not a token, as its impossible to learn correlations with it. Were one to introduce a true space token, it would be interpretted by the retokenizer as being an invisible letter, and learned patterns would no longer correspond to linguistic units. It would instead learn an interpretation of the data reflecting artifacts of the tokenization.
Thus we present the model here using a basis tokenset of lowercase alphabetic characters. The dimension of the alphabet can be anything so long as the tokenset is intrinsically meaningless. Not because we necessarily need meaningless tokens, but only because the retokenizer treats all tokens as such regardless.
Boundaries are statistical & emergent. This model does not make use of explicit boundary marking. Hierarchical boundaries, at any level of language structure, are an implicit quality of the model. This is discussed more in the context of syntax in Sec 8.
Figure 3. Comparing Hamiltonians of topologically protected global order vs unprotected. Context overlapping with a feature is driven to the fixed point (energy extremum) by a restoring force. We use convention for energy E=HE=-H.
Figure 4. Hierarchical short-term memory retrieves via flows toward terminal nodes. Multiple terminal nodes compete during retrieval if they share features. However, retrieval becomes scale invariant beyond last overlap. Note: DAG’s are disentangled during formation of long-term memory—see Fig 7.
Figure 5. 500 words from Alice & Wonderland (GutenbergAlice). Solid line is log-normal fit. The shortest lengthscale patterns are biased by the mechanical constraints of pronunciation/listening. But these biases don’t account for why word length is finite. Structural constraints of memory become relevant at length scales $n>n_{\text{peak}}$, leading to the $d_{n}$ collapse, and limiting the length of the longest pattern.
4. Universal Sub-Word Structure
Here we strip away all word meaning and focus on the statistical structure of the underlying patterns.
This structure is quantified in terms of two distributions: (1) the number of unique $n$-grams (a.k.a word-forming patterns) $d_{n}$ versus length $n$; and (2) the frequency of $n$-grams rank-ordered from most to least frequent (w/ one plot per $n$).
4.1. Words as finite patterns & $d_{n}$
The quantity $d_{n}$ measures the width of the feature space, and is grown layer-by-layer in the model. It has been known for at least half a century that $d_{n}$ approximates a log-normal distribution for both subword patterns (Herdan1958_logNormal_intraword) and syntax (Williams1940_logNormal_interwords).
Such a distribution (Fig 5) rises exponentially to a smooth peak at $n_{\text{peak}}$ before collapsing to zero at $n_{\text{max}}$. This collapse to zero coincides with the reality of finite word length, and how it scales with $n$ hints at the structural constraints that enforce that reality.
This distribution differs from $d_{n}$ for unstructured noise (pulled from a uniform distribution)—see Fig 11 in Sec A.2. For noise w/o spaces, $n_{\text{max}}$ is equal to the length of the spaceless string, which is arbitrary. Smaller chunks can be formed by randomly inserting spaces at some frequency, which places a cutoff on length; but the transition across $n_{\text{peak}}$ is sharp (not log-normal), owing to the fact that $n_{\text{peak}} & n>n$$n_{\text{peak}}$ are governed by different distributions. Left of the peak grows exponentially as $d^{n}$, because there is sufficient random statistics to guarantee the smaller $n$-grams are each observed at least once.
Note however that noise, while unstructured, has structure living inside of it. As we pointed out above, every configuration will be observed given enough noise. By constructing an appropriate filter, we can sift out a structured distribution. One would then describe the filter (not the noise) as being the source of the structured data. The noise is just the catalyst for learning off the filter.
The retokenizer does just this. We have a retokenizer self-organize, seeded by the same spaceless noise string described above (Fig 11). We plot the memories of the retokenizer alongside both the spaceless training string and its random chunks. It’s peak is smooth and fits a log-normal.
Note that one does not need fancy tensor calculations to reproduce this result, but can do this themselves by hand using the very same noise data available in Sec A.2. The rules: (1) If a bigram shows up at least twice, put it in the graph. (2) For $n>2$, keep all $n$-grams which are smooth mergers of the relevant bigrams. (3) The accepted strings are the terminal nodes of the graph.
No arbitrary cutoff is imposed beyond $n=2$, as the smoothness constraint is strong enough to drive the collapse.
One doesn’t need to explicitly generate noise to learn from in order to generate a DAG graph, but can grow a random DAG from a random subset of bigrams and their mergers (Sec A.3). The smoothness acts as a hierarchical constraint, which enforces closure over all possible growths. (We explore these closed hierarchical structures in the next section.)
In formal language terms, the subregular symbolic structure generates two types of strings: (1) finite strings living at terminal nodes; and (2) a small subset of simple repeating infinite strings, which arise due to self-mergers (e.g $aa\cdot a\rightarrow aaa$). Such self-mergers are more likely to observe with random growths for small $d$; and for any $d$, only extend to infinite in the special limit [^11] It’s worth pointing out that this makes them “fine-tuned” in the sense that they aren’t physically expected. However, there are some simple self-mergers which show up in writing, e.g from lol: lolololol w/ various length. where no additional cutoff is imposed for $n>2$. And self-organizing against a noise string makes them exponentially unlikely, because they have to be observed in the string.
4.2. Hierarchical Feature Mixing & Random Language
We begin by analyzing the structure of a randomly sampled vocabulary, consisting of 100 English or Spanish personal names (in Sec A.1). We plot the $d_{n}$ distribution for this set, which follows an approximately log-normal form, with the largest $n$ determined by the longest name in the sample.
In contrast, cutting closer to $n_{\text{peak}}$ loosens the scale-dependent constraints, allowing novel $n$-grams to form. Nevertheless, even here, the surviving short-scale DAG structure continues to guide regrowth, guaranteeing closure.
We find that though the original notion of smoothness is perfectly preserved below the cut, cutting away more of the $d_{n}$ peak causes the hierarchy to begin losing its recognizable English/Spanish nature. This highlights the importance of the $n_{\text{peak}}$ grams to language structure [^12] For $n\lesssim n_{\text{peak}}$, $d_{n}$ grows approximately exponentially, with learned constraints ($g^{(n)}=0$) pruning branches and thereby reducing the effective growth rate. For $n>n_{\text{peak}}$, the smoothness condition takes over: new tokens are permitted only if all subgrams are valid, which recursively restricts composition and causes a collapse in $d_{n}$. This collapse guarantees finite word length. In formal language terms, this might be described as a consequence of the subregular limitations of stored strings..
Finally, if we cut at $n_{\text{cut}}=2$, we discard all memory of the original names. This is the data-free regime, where growing a hierarchy generates what we refer to as a “random language” (Eugenio2025_HebbianLanguage). Nevertheless, growing hierarchies under the smoothness constraint still yields finite word-lengths and an approximately log-normal $d_{n}$ distribution (Herdan1958_logNormal_intraword; Williams1940_logNormal_interwords; Eckhard&Werner&Markus2001_logNormal_review). This shows that bounded, hierarchical morphology is an intrinsic feature of the model itself, independent of any training data, and that language structure arises naturally from smooth hierarchical learning.
4.3. Synaptic Zipf-laws
The rank-ordered frequencies (a.k.a Zipfian plots (Zipf1949; Zanette&Montemurro2002)) measure the relevancy (ranked from most to least frequent) of those features in the sample—Fig 6. We can interpret these ranked frequencies as defining an effective memory spectra $g^{(n)}_{\text{eff}}$, where Hebbian learning imprints frequency counts from local events $\mathcal{E}_{n}$ directly into memory:
$g^{(n)}_{\text{eff}}=\sum_{\mathcal{E}_{n}^{\small\text{vocab}}}\tilde{v}v,$
with one spectra per level $n$.
Random languages generated by our model exhibit rank-ordered $n$-gram frequency distributions that follow a scale-invariant Zipf-like law (Gan&Wang&Han2009_NtupleZipf).
At short lengths ($n<n_{\text{peak}}$), constraints are minimal, and the distribution appears structureless: a Zipf-like power-law head followed by a tail of irrelevant $n$-grams, similar to those found in uniform random strings (Eugenio2025_HebbianLanguage; Gan&Wang&Han2009_NtupleZipf)—Fig 13. However, as $n$ increases, the DAG constraint increasingly shapes memory formation. Beyond the peak, the tail vanishes and the power-law scaling persists, signaling the emergence of scale-invariance tied to a dominant compositional structure at longer lengths.
4.3.1. Scale-Invariance
Note that the widest spectrum of Fig 6 is $n=n_{\text{peak}}$ due to the rank ordering (largest dimension). The earlier layers (left of the $d_{n}$ peak) arise out of pure randomness, and initially grows exponentially. The DAG constraint grows stronger with increasing $n$, until $n_{\text{peak}}$, beyond which it becomes the dominate effect.
This gives rise to a characteristic length scale $l_{\text{token}}$, which we here dub the invariant token length: Measures the distance past $n_{\text{peak}}$ to the last DAG overlap (where slope vanishes). Beyond which all autoregressive flows are guaranteed to converge to a single terminal node.
Note that $l_{\text{token}}\simeq 0$ for noise. This is to say that that the more noisy the string, the less likely its overlap with other memories, and therefore the less its interference during retrieval from short-term memory. Memorizing noise requires learning all $n$-grams (so keep all bigrams), and can only be done for a limited length string, due to the unchecked scaling in the feature dimension. Finite $l_{\text{token}}$ indicates local structure, with a larger number of DAG overlaps, and hence requires lower overhead for a DAG memory.
Note the distinction between scale invariance of a retrieving memory, and the scale invariance of composing subword patterns (the persistence of the power-law slopes). The latter arises from contributions from overlapping DAG graphs, whose slopes decrease as we approach longer length scales & fewer overlaps. Once these slopes vanish, a single memory is guaranteed for a pattern of that length. Hence the vanishing of the slopes indicates increasing scale invariance during retrieval.
[…]
4.3.2. Local Ordering, the Illusion of Utility, & Inter-Generational Learning
Figure 14 shows the rank-ordered frequency spectra for Greek words in the Rosetta Stone. Removing spaces (Fig. 14-b) does not destroy the compositional structure: the invariant token length $l_{\text{token}}$ remains finite. This indicates that word-bearing strings are locally ordered: a restricted subset of short-range patterns dominates the statistics even when boundaries are erased. In other words, the local structure survives boundary removal, suggesting that underlying words can in principle be inferred from undifferentiated spaceless text.
This possibility motivates “token-finding” approaches, where words are inferred directly from statistical regularities. Such methods have utility for parsing undeciphered strings or segmenting continuous speech. And it has been suggested that the lack of explicit boundary markers in human speech suggest children may exploit this during word learning (Brent1999_SpeechSegmentation&WordDiscovery). Modern statistical approaches similarly attempt unsupervised token discovery (Anonymous2024_LatentSegmentLanguageModels; Pagnoni2024_ByteLatentTransformerPatches; Schmidt2025_BoundlessBytePairEncoding; Goldwater&Griffiths&Johnson2009_BayesianWordSegmentation; Creutz&Lagus2006_UnsupervisedMorphologyInduction_Morfessor).
However, this behavior is not a property of the retokenizer. The model cannot, in general, recover the “correct” words from a spaceless string. While it can form symbolic tokens from unsegmented input, those tokens must first be demonstrated—that is, the model must learn what constitutes a “good” pattern. Because it operates purely on local correlations, it is sensitive only to the frequency statistics of nearby patterns. The same local order can be mimicked by copy-&-paste methods (Simon-like processes) for generating language-like strings (Gan&Wang&Han2009_NtupleZipf). There are no actual words composing those strings, but have a compositional structure due to copy-&-pasting string segments to the end of the string.
If trained on such sequences, the retokenizer would simply invent its own symbols consistent with its inductive biases, just as it does when trained on noise. Even on real text containing concatenation errors, it may merge frequent sequences into composite tokens (e.g learning theboy instead of the and boy).
Crucially, this local ordering is not a designed feature of human language, nor an adaptation to aid learners. It is a consequence of locality—a universal property of any microscopic system composed of locally interacting parts. Any utilitarian structure is purely coincidental. Children do not learn “correct” words but form internal interpretations aligned with their own inductive biases. If those biases resemble those of the retokenizer, then even “incorrect” segmentations still preserve the universal subword statistics. Universal substructure is a dynamical fixed point of local learning; it cannot collapse under continued relearning (Shumailov&Gal2024nature_ModelCollapse; Shumailov2024_CurseOfRecursion). Thus, linguistic change can be viewed as the reinterpretation of learned errors, continuously recycled into good language patterns by the same universal dynamics.
Figure 6. The frequency an $n$-gram (length $n$) appears in the vocabulary, rank-ordered from most to least frequent. The vocabulary is for the randomly grown language bjfyi (in supplement). Equivalently, this plot can be understood as representing an effective model spectra ($g^{(n)}_{\text{eff}}$ per $n$). Scale-invariance emerges beyond $n_{\text{peak}}=3$ due to the compositional structure.
@@ 5. The Benefits of Thinking About It
The retokenizer is not a very good model. It’s bifix-free subregular language structure limits the complexity of strings it can represent. Nor is possible to repurpose it as a tokenizer (pre-processor) for an LLM. This would amount to discarding its hierarchical structure and collapsing all learned $n$-grams into a flat static vocabulary, which would scale poorly and break the model’s inductive biases.
Its value is in its ability to explain where basic word structure comes from, and its relationship to symbolic order. Its word symbols are entirely meaningless, which teaches that us universal structure is independent to meaning. But this explanation is obviously incomplete.
So then: “What is microscopically possible?” As a general statement, local memories are simple retrieval memories which capture the trace of an event. They aren’t that coordinated, and can only do two things: retrieve or forget. Therefore, continued learning arises from the retrieval.
@@ 5.1. The Replay Game (Model Thinking)
$\text{e.g}\;\;\;a_{\alpha={\bf cats}}\leftrightarrow\left\{v_{\bf c},v_{\bf a},v_{\bf t},v_{\bf s};\tilde{v}^{(2)}_{\bf ca},\tilde{v}^{(2)}_{\bf at},\tilde{v}^{(2)}_{\bf ts};\tilde{v}^{(3)}_{\bf cat},\tilde{v}^{(3)}_{\bf ats};\tilde{v}^{(4)}_{\bf cats}\right\}$.
The result is a key-value memory binding (Tyulmankov&all2021_BioKeyValue; Gersham&Fiete&Irie2025_KeyValueMemory; Ba&all2016_FastWeightsAssociativeKeyValue): the neuron $a_{\alpha}$ becomes a long-term embedding associated with the full replayed word. Over time, this mechanism constructs a library of long-term memory neurons $\{a_{\alpha}\}$, each associated with a unique word and its compositional features.
5.1.1. Non-static architecture, protected retrieval
Long-term memories live in parallel channels; adding one does not overwrite others. It is a content-addressable memory, sharing qualities with sparse distributed memories (Kanerva1988_SDM) & dynamic associative networks (Beavers2010_DAN1; Beavers&Harrison_DAN2). Memory capacity is infinite, as tokens can be stored onto the disk until needed. Nothing is ever overwritten, and thus no catastrophic forgetting.
The parallelization provides topological protection (replay or recognition of one memory does not interfere with another). Thus can retrieve & accept hyper independent from hyperbolic.
Because the memory is local, each token decomposes into a series of small tensor contractions. This supports staged (& partial [^13] Partial-key checks: You don’t need to compute the whole token contribution, only up to the part where you can be certain no other token is more similar. This is similar to how some subword patterns like ndwi necessarily only converge to sandwich—a length beyond which retrieval is scale-invariant–Fig 4.) key checks for efficient lookup on serial hardware, while an ideal neuromorphic device parallelizes the entire search.
5.1.2. Sparse not dense
Our model is actually sparse—we just use a convenient dense representation. This produces redundancy in the form of gauges, which we exploit to perform right & left token predictions. Basis tokens are stored locally as one-hot encodings, not random memory vectors. Structure emerges from local projector compression (a consequence of locality); and the distributed nature of the memory post-replay arises due to simultaneity of those replayed features. Thus structure can be learned without any data [^14] Training a model, even a local model, with explicit language data is technically a non-local process. The specific ordered strings of that data were prepared by a non-local source—the human modeler. It would not have learned the word “cat” had you not ordered the data. or human intervention.
This is completely unlike dense distributed memories, which typically assume a basis of random memory vectors in order to reduce overlap (Hopfield1982; HopfieldAllNeed_Ramsauer2021; LongSequenceHopfield_Hamza2023). There are no binding/unbinding operations as in Smolensky’s tensor product formulation (Smolensky1990).
5.1.3. Topology before meaning
The retokenizer operates over meaningless topological symbols (bifix-free, subregular). This yields exact, boundary-aware retrieval paths: a terminal node is reachable only through its smooth subpaths, so inferred boundaries are statistical fixed points of the local dynamics rather than externally imposed marks. This guarantees symmetry between morphology and syntax—both arising from correlations encoding subregular strings of the retokenizer. The tree structure of the embedding graphs are identical to syntax trees.